Shelly has a container of colored tiles: 3 green, 4 blue, 6 yellow, and 12 red. The first tile she takes from the container is red. If she does not put the red tile back into the container, what is the probability that the second tile she randomly takes out of the container will be red?

1
Expert's answer
2021-05-18T14:13:56-0400

"P=\frac{11}{3+4+6+12-1}=\frac{11}{24}"